Test at the α = .01 significance level whether the mean of a random sample of size n = 16 is statistically significantly less than 10. If the distribution from which the sample was taken is normal, x¯ = 8.4 and σ2 = 10.24. What are the null and alternative hypotheses for this test?



1
Expert's answer
2021-05-25T15:12:37-0400

Hypothesized Population Mean "\mu=10"

Population Standard Deviation "\sigma=\sqrt{\sigma^2}=\sqrt{10.24}=3.2"

Sample Size "n=16"

Sample Mean "\bar{x}=8.4"

Significance Level "\alpha=0.01"


Null and Alternative Hypotheses

The following null and alternative hypotheses need to be tested:

"H_0: \mu\geq10"

"H_1: \mu<10"

This corresponds to a left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.01" and the critical value for  left-tailed test is "z_c=-2.3263." 

The rejection region for this left-tailed test is "R=\{z:z<-2.3263\}."


The "z" - statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{8.4-10}{3.2/\sqrt{16}}=-2"

Since it is observed that "z=-2>-2.3263=z_c," it is then concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is less than "10," at the "\alpha=0.01" significance level.


Using the P-value approach: The p-value for one-tailed, the significance level "\alpha=0.01, z=-2, \text{left-tailed}" is "p=0.02275," and since "p=0.02275>0.01=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu" is less than "10," at the "\alpha=0.01" significance level.